### v2

@@ 5.2 Sparsity enhanced low-rank tensor regularization
Taking $\mathbf{R}_{nl}$ as an example, in the $\mathbf{R}_{nl}(\mathcal{L})$ that we construct, the first dimension indicates the spatial information, the second dimension reveals the nonlocal self-similarity, the third dimension reflects the spectral correlation, and the fourth dimension shows no correlation. Hence, we adopt the independent 3-D HOSVD to obtain a low-rank approximation of $\mathbf{R}(\mathcal{L})$, that is,
$\mathbf{R}(\mathcal{L})\approx\mathcal{[}\mathcal{G}]\times_{1}[{X}_{1}]\times_{2}[{X}_{2}]\times_{3}[{X}_{3}],$
where $[\mathcal{G}]\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}\times N}$ denotes independent core tensors, and $[{X}_{i}]\in\mathbb{R}^{m_{i}\times n_{i}\times N}$ denotes the $i$-th factor matrices such that $[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}]$.
To further boost the low-rankness of $\mathbf{R}(\mathcal{L})$, we propose a sparsity-enhanced nonlocal low-rank tensor regularization term as follows
(25) $\frac{1}{2}\|\mathcal{[}\mathcal{G}]\times_{1}[{X}_{1}]\times_{2}[{X}_{2}]\times_{3}[{X}_{3}]-\mathbf{R}(\mathcal{L})\|^{2}_{F}+\|[\mathcal{G}]\|_{1,w},$
where $\|[\mathcal{G}]\|_{1,w}$ is given in (3) with $w\in\mathbb{R}^{N}_{++}$. In particular, the first term of (25) measures the closeness between $\mathbf{R}(\mathcal{L})$ and the approximated low-rank tensor, and the second term measures the sparsity of the independent core tensors $[\mathcal{G}]$. Using the proximal operator of the $\ell_{1}$ norm, the proximal operator of $\|\cdot\|_{1,w}$ can be computed component-wisely by the soft thresholding operator as follows
$\left(\operatorname{prox}_{\|\cdot\|_{1,w}}([\mathcal{G}])\right)_{i_{1}i_{2}i_{3}}^{(j)}$ $=\operatorname{prox}_{w_{j}|\cdot|}\left([\mathcal{G}]^{(j)}_{i_{1}i_{2}i_{3}}\right)$
@@ 5.3 Tensor $\ell_{2,p}$ norm for group sparsity
(27) $\min_{s}\mu\|s\|^{p}_{2}+\frac{1}{2}\|s-\tilde{s}\|^{2}_{2},$
where $s\in\mathbb{R}^{I_{1}}$ and $\tilde{s}\in\mathbb{R}^{I_{1}}$, for simplicity, represent $s_{:i_{2}i_{3}}$ and $\tilde{s}_{:i_{2}i_{3}}$, respectively, and $\|s\|^{p}_{2}=(s_{1}^{2}+s_{2}^{2}+\cdots+s_{I_{1}}^{2})^{\frac{p}{2}}$. It follows from the triangle inequality that the objective function of (27) satisfies the following inequality for any $s\in\mathbb{R}^{I_{1}}$
(28) $\mu\|s\|^{p}_{2}+\frac{1}{2}\|s-\tilde{s}\|^{2}_{2}\geq\mu\|s\|^{p}_{2}+\frac{1}{2}(\|s\|_{2}-\|\tilde{s}\|_{2})^{2}.$
And the equality holds if and only if $s=t\tilde{s}$ for some $t\geq 0$ or $\tilde{s}=0$. Observe that the right-hand side of the inequality is only related to $\|s\|_{2}$ and $\|\tilde{s}\|_{2}$. If $\tilde{s}=0$, the solution of problem (27) is $s=0$. If $\tilde{s}\neq 0$, we can view $s$ as $s=t\|\tilde{s}\|_{2}v$ with $t\geq 0$ being a scalar and $v\in\mathbb{R}^{I_{1}}$ being a unit vector. When we restrict the minimization problem (27) by $\|s\|_{2}=t\|\tilde{s}\|_{2}$ with a fixed $t$, according to (28), the solution of the restricted problem of (27) is obtained only when $v=\frac{\tilde{s}}{\|\tilde{s}\|_{2}}$. Hence, if $\tilde{s}\neq 0$, the solution of (27) is $s=t\tilde{s}$, where $t$ is a minimizer of the following problem
@@
with $\nu=\mu\|\tilde{s}\|_{2}^{p-2}$. That is, it only requires to solve a one-dimensional problem (29) for computing the solutions of problem (27).
@@
$\Gamma_{\mu}(\|\tilde{s}_{:i_{2}i_{3}}\|_{2})\tilde{s}_{:i_{2}i_{3}}\in\operatorname{prox}_{\mu\|\cdot\|_{2}^{p}}(\tilde{s}_{:i_{2}i_{3}})=\left(\operatorname{prox}_{\mu\|\cdot\|_{2,p}^{p}}(\widetilde{\mathcal{S}})\right)_{:i_{2}i_{3}},$
@@
$s_{:i_{2}i_{3}}^{k+1}=\Gamma_{\tilde{\gamma}}\left(\left\|\tilde{s}^{k}_{:i_{2}i_{3}}\right\|_{2}\right)\tilde{s}^{k}_{:i_{2}i_{3}},$
where $\tilde{\mathcal{S}}^{k}=\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D}\right)$, and $\tilde{s}^{k}_{:i_{2}i_{3}}$ is the $(i_{2},i_{3})$-th mode-$1$ fibers of $\tilde{\mathcal{S}}^{k}$.
@@ 5.4 Two-phase MLTL2p method
The spectral correlation and spatial nonlocal self-similarity in HSIs imply multi-scale low-rankness, which includes global low-rankness, local block low-rankness, and nonlocal self-similar tensor low-rankness. In the following, we propose a multi-scale low-rank model that imposes low-rankness on global, local, and nonlocal tensors extracted via $\mathbf{R}_{g}$, $\mathbf{R}_{l}$, and $\mathbf{R}_{nl}$, respectively.
(31) $\begin{split}\min_{\begin{subarray}{c}[{X}_{i}]_{g},[{X}_{i}]_{l},[{X}_{i}]_{nl},\\ [\mathcal{G}]_{g},[\mathcal{G}]_{l},[\mathcal{G}]_{nl},\mathcal{S},\mathcal{L}\end{subarray}}\quad&\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\|\mathcal{S}\|^{p}_{2,p}\\ &+\|[\mathcal{G}]_{g}\|_{1,w}+\frac{\delta_{g}}{2}\|\mathbf{R}_{g}(\mathcal{L})-[\mathcal{G}]_{g}\times_{1}[X_{1}]_{g}\times_{2}[X_{2}]_{g}\times_{3}[X_{3}]_{g}\|^{2}_{F}\\ &+\|[\mathcal{G}]_{l}\|_{1,w}+\frac{\delta_{l}}{2}\|\mathbf{R}_{l}(\mathcal{L})-[\mathcal{G}]_{l}\times_{1}[X_{1}]_{l}\times_{2}[X_{2}]_{l}\times_{3}[X_{3}]_{l}\|^{2}_{F}\\ &+\|[\mathcal{G}]_{nl}\|_{1,w}+\frac{\delta_{nl}}{2}\|\mathbf{R}_{nl}(\mathcal{L})-[\mathcal{G}]_{nl}\times_{1}[X_{1}]_{nl}\times_{2}[X_{2}]_{nl}\times_{3}[X_{3}]_{nl}\|^{2}_{F}\\ \mbox{s.t. }&[{X}_{i}]_{g}^{\top}[{X}_{i}]_{g}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\\ &[{X}_{i}]_{l}^{\top}[{X}_{i}]_{l}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\\ &[{X}_{i}]_{nl}^{\top}[{X}_{i}]_{nl}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3.\end{split}$
@@
1:  Initialize $(\mathcal{S}^{0},[{X}_{i}^{0}]_{g},[{X}_{i}^{0}]_{l},[{X}_{i}^{0}]_{nl},[\mathcal{G}^{0}]_{g},[\mathcal{G}^{0}]_{l},[\mathcal{G}^{0}]_{nl},\mathcal{L}^{0})$ with $[{X}_{i}^{0}]_{g},[{X}_{i}^{0}]_{l},[{X}_{i}^{0}]_{nl}$ being independently orthogonal;
2:  Set the tensor extraction operators $\mathbf{R}_{g}$, $\mathbf{R}_{l}$, and $\mathbf{R}_{nl}$;
3:  Set parameters $\alpha_{\mathcal{S}},\alpha_{X},\alpha_{\mathcal{G}}>0$;
@@
$\mathcal{L}^{k+1}=$ $\left(\delta_{g}\mathcal{W}_{\mathbf{R}_{g}}+\delta_{l}\mathcal{W}_{\mathbf{R}_{l}}+\delta_{nl}\mathcal{W}_{\mathbf{R}_{nl}}+\operatorname{Id}\right)^{-1}$
$\odot\Big{(}\delta_{g}\mathbf{R}_{g}^{\top}([\mathcal{Y}^{k+1}]_{g})+\delta_{l}\mathbf{R}_{l}^{\top}([\mathcal{Y}^{k+1}]_{l})$
$+\delta_{nl}\mathbf{R}_{nl}^{\top}([\mathcal{Y}^{k+1}]_{nl})+(\mathcal{D}-\mathcal{S}^{k+1})\Big{)};$
@@
11:  $(\mathcal{S}^{k},[{X}_{i}^{k}]_{g},[{X}_{i}^{l}]_{l},[{X}_{i}^{k}]_{nl},[\mathcal{G}^{k}]_{g},[\mathcal{G}^{k}]_{l},[\mathcal{G}^{k}]_{nl},\mathcal{L}^{k})$
To enhance the efficiency and effectiveness of the MLTL2p method, we implement a two-phase scheme on the multi-scale low-rank model and the P-BCD method. In the first phase, we employ only the global and local low-rank regularization, which is sufficient for identifying stripes and dead lines. In particular, we use Algorithm 2 to solve model (31) with the parameter $\delta_{nl}$ set to zero. In the second phase, we utilize the nonlocal low-rank regularization along with the global and local low-rank regularization. As the estimated HSI from the first phase has already removed most of the noise, we can perform only one block matching to obtain a fixed nonlocal block extraction $\mathbf{R}_{nl}$, which can significantly reduce the computational cost. By applying Algorithm 2 to solve model (31) in this phase, the algorithm can refine additional details in the HSI, thereby improving the quality of the restored HSI.
@@ 6 Experimental Results
In this section, we conduct numerical experiments for removing mixed noise in HSIs. We compare the proposed methods with five methods, which are BM4D [28] for removing Gaussian noise, and LRTD [12], SNLRSF [6], LRTFL0 [40] and QRNN3D [38] for removing mixed noise. All numerical experiments are implemented in Matlab R2018a and executed on a personal desktop (Intel Core i7 9750H at 2.60 GHz with 16 GB RAM).

### v3

@@ 5.2 Sparsity enhanced low-rank tensor regularization
Taking $\mathscr{R}_{nl}$ as an example, in the $\mathscr{R}_{nl}(\mathcal{L})$ that we construct, the first dimension indicates the spatial information, the second dimension reveals the nonlocal self-similarity, the third dimension reflects the spectral correlation, and the fourth dimension shows no correlation. Hence, we adopt the independent 3-D HOSVD to obtain a low-rank approximation of $\mathscr{R}(\mathcal{L})$, that is,
$\mathscr{R}(\mathcal{L})\approx\mathcal{[}\mathcal{G}]\times_{1}[\boldsymbol{X}_{1}]\times_{2}[\boldsymbol{X}_{2}]\times_{3}[\boldsymbol{X}_{3}],$
where $[\mathcal{G}]\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}\times N}$ denotes independent core tensors, and $[\boldsymbol{X}_{i}]\in\mathbb{R}^{m_{i}\times n_{i}\times N}$ denotes the $i$-th factor matrices such that $[\boldsymbol{X}_{i}]^{\top}[\boldsymbol{X}_{i}]=[\boldsymbol{I}_{n_{i}}]$.
To further boost the low-rankness of $\mathscr{R}(\mathcal{L})$, in additional to restricting the core tensors to a unified size, we propose a sparsity-enhanced nonlocal low-rank tensor regularization term as follows
(26) $\frac{1}{2}\|\mathcal{[}\mathcal{G}]\times_{1}[\boldsymbol{X}_{1}]\times_{2}[\boldsymbol{X}_{2}]\times_{3}[\boldsymbol{X}_{3}]-\mathscr{R}(\mathcal{L})\|^{2}_{F}+\|[\mathcal{G}]\|_{1,\boldsymbol{w}},$
where $\|[\mathcal{G}]\|_{1,\boldsymbol{w}}$ is given in (3) with $w\in\mathbb{R}^{N}_{++}$. In particular, the first term of (26) measures the closeness between $\mathscr{R}(\mathcal{L})$ and the approximated low-rank tensor, and the second term measures the sparsity of the independent core tensors $[\mathcal{G}]$. Using the proximal operator of the $\ell_{1}$ norm, the proximal operator of $\|\cdot\|_{1,\boldsymbol{w}}$ can be computed component-wisely by the soft thresholding operator as follows
$\left(\operatorname{prox}_{\|\cdot\|_{1,\boldsymbol{w}}}([\mathcal{G}])\right)_{i_{1}i_{2}i_{3}}^{(j)}$ $=\operatorname{prox}_{w_{j}|\cdot|}\left([\mathcal{G}]^{(j)}_{i_{1}i_{2}i_{3}}\right)$
@@ 5.3 Tensor $\ell_{2,p}$ norm for group sparsity
(28) $\min_{\boldsymbol{s}}\mu\|\boldsymbol{s}\|^{p}_{2}+\frac{1}{2}\|\boldsymbol{s}-\tilde{\boldsymbol{s}}\|^{2}_{2},$
where $\boldsymbol{s}\in\mathbb{R}^{I_{1}}$ and $\tilde{\boldsymbol{s}}\in\mathbb{R}^{I_{1}}$, for simplicity, represent $\boldsymbol{s}_{:i_{2}i_{3}}$ and $\tilde{\boldsymbol{s}}_{:i_{2}i_{3}}$, respectively, and $\|\boldsymbol{s}\|^{p}_{2}=({s}_{1}^{2}+{s}_{2}^{2}+\cdots+{s}_{I_{1}}^{2})^{\frac{p}{2}}$. It follows from the triangle inequality that the objective function of (28) satisfies the following inequality for any $s\in\mathbb{R}^{I_{1}}$
(29) $\mu\|\boldsymbol{s}\|^{p}_{2}+\frac{1}{2}\|\boldsymbol{s}-\tilde{\boldsymbol{s}}\|^{2}_{2}\geq\mu\|\boldsymbol{s}\|^{p}_{2}+\frac{1}{2}(\|\boldsymbol{s}\|_{2}-\|\tilde{\boldsymbol{s}}\|_{2})^{2}.$
And the equality holds if and only if $\boldsymbol{s}=t\tilde{\boldsymbol{s}}$ for some $t\geq 0$ or $\tilde{\boldsymbol{s}}=0$. Observe that the right-hand side of the inequality is only related to $\|\boldsymbol{s}\|_{2}$ and $\|\tilde{\boldsymbol{s}}\|_{2}$. If $\tilde{\boldsymbol{s}}=0$, the solution of problem (28) is $\boldsymbol{s}=\boldsymbol{0}$. If $\tilde{\boldsymbol{s}}\neq\boldsymbol{0}$, we can view $s$ as $s=t\|\tilde{\boldsymbol{s}}\|_{2}\boldsymbol{v}$ with $t\geq 0$ being a scalar and $v\in\mathbb{R}^{I_{1}}$ being a unit vector. When we restrict the minimization problem (28) by $\|\boldsymbol{s}\|_{2}=t\|\tilde{\boldsymbol{s}}\|_{2}$ with a fixed $t$, according to (29), the solution of the restricted problem of (28) is obtained only when $\boldsymbol{v}=\frac{\tilde{\boldsymbol{s}}}{\|\tilde{\boldsymbol{s}}\|_{2}}$. Hence, if $\tilde{\boldsymbol{s}}\neq\boldsymbol{0}$, the solution of (28) is $\boldsymbol{s}=t\tilde{\boldsymbol{s}}$, where $t$ is a minimizer of the following problem
@@
with $\nu=\mu\|\tilde{\boldsymbol{s}}\|_{2}^{p-2}$. That is, it only requires to solve a one-dimensional problem (30) for computing the solutions of problem (28).
@@
$\Gamma_{\mu}(\|\tilde{\boldsymbol{s}}_{:i_{2}i_{3}}\|_{2})\tilde{\boldsymbol{s}}_{:i_{2}i_{3}}\in\operatorname{prox}_{\mu\|\cdot\|_{2}^{p}}(\tilde{\boldsymbol{s}}_{:i_{2}i_{3}})=\left(\operatorname{prox}_{\mu\|\cdot\|_{2,p}^{p}}(\widetilde{\mathcal{S}})\right)_{:i_{2}i_{3}},$
@@
$\boldsymbol{s}_{:i_{2}i_{3}}^{k+1}=\Gamma_{\tilde{\gamma}}\left(\left\|\tilde{\boldsymbol{s}}^{k}_{:i_{2}i_{3}}\right\|_{2}\right)\tilde{\boldsymbol{s}}^{k}_{:i_{2}i_{3}},$
where $\tilde{\mathcal{S}}^{k}=\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D}\right)$, and $\tilde{\boldsymbol{s}}^{k}_{:i_{2}i_{3}}$ is the $(i_{2},i_{3})$-th mode-$1$ fibers of $\tilde{\mathcal{S}}^{k}$.
@@ 5.4 Two-phase MLTL2p method
The spectral correlation and spatial nonlocal self-similarity in HSIs imply multi-scale low-rankness, which includes global low-rankness, local block low-rankness, and nonlocal self-similar tensor low-rankness. In the following, we propose a multi-scale low-rank model that imposes low-rankness on global, local, and nonlocal tensors extracted via $\mathscr{R}_{g}$, $\mathscr{R}_{l}$, and $\mathscr{R}_{nl}$, respectively.
(32) $\begin{split}\min_{\begin{subarray}{c}[\boldsymbol{X}_{i}]_{g},[\boldsymbol{X}_{i}]_{l},[\boldsymbol{X}_{i}]_{nl},\\ [\mathcal{G}]_{g},[\mathcal{G}]_{l},[\mathcal{G}]_{nl},\mathcal{S},\mathcal{L}\end{subarray}}\quad&\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\|\mathcal{S}\|^{p}_{2,p}\\ +&\|[\mathcal{G}]_{g}\|_{1,\boldsymbol{w}}+\frac{\delta_{g}}{2}\|\mathscr{R}_{g}(\mathcal{L})-[\mathcal{G}]_{g}\times_{1}[\boldsymbol{X}_{1}]_{g}\times_{2}[\boldsymbol{X}_{2}]_{g}\times_{3}[\boldsymbol{X}_{3}]_{g}\|^{2}_{F}\\ +&\|[\mathcal{G}]_{l}\|_{1,\boldsymbol{w}}+\frac{\delta_{l}}{2}\|\mathscr{R}_{l}(\mathcal{L})-[\mathcal{G}]_{l}\times_{1}[\boldsymbol{X}_{1}]_{l}\times_{2}[\boldsymbol{X}_{2}]_{l}\times_{3}[\boldsymbol{X}_{3}]_{l}\|^{2}_{F}\\ +&\|[\mathcal{G}]_{nl}\|_{1,\boldsymbol{w}}+\frac{\delta_{nl}}{2}\|\mathscr{R}_{nl}(\mathcal{L})-[\mathcal{G}]_{nl}\times_{1}[\boldsymbol{X}_{1}]_{nl}\times_{2}[\boldsymbol{X}_{2}]_{nl}\times_{3}[\boldsymbol{X}_{3}]_{nl}\|^{2}_{F}\\ \mbox{s.t. }~{}~{}&[\boldsymbol{X}_{i}]_{g}^{\top}[\boldsymbol{X}_{i}]_{g}=[\boldsymbol{I}_{n_{i}}],\quad i=1,2,3,\\ &[\boldsymbol{X}_{i}]_{l}^{\top}[\boldsymbol{X}_{i}]_{l}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\\ &[\boldsymbol{X}_{i}]_{nl}^{\top}[\boldsymbol{X}_{i}]_{nl}=[\boldsymbol{I}_{n_{i}}],\quad i=1,2,3.\end{split}$
@@
1:  Initialize $(\mathcal{S}^{0},[\boldsymbol{X}_{i}^{0}]_{g},[\boldsymbol{X}_{i}^{0}]_{l},[\boldsymbol{X}_{i}^{0}]_{nl},[\mathcal{G}^{0}]_{g},[\mathcal{G}^{0}]_{l},[\mathcal{G}^{0}]_{nl},\mathcal{L}^{0})$ with $[\boldsymbol{X}_{i}^{0}]_{g},[\boldsymbol{X}_{i}^{0}]_{l},[\boldsymbol{X}_{i}^{0}]_{nl}$ being independently orthogonal;
2:  Set the tensor extraction operators $\mathscr{R}_{g}$, $\mathscr{R}_{l}$, and $\mathscr{R}_{nl}$;
3:  Set parameters $\alpha_{\mathcal{S}},\alpha_{\boldsymbol{X}},\alpha_{\mathcal{G}}>0$;
@@
$\mathcal{L}^{k+1}=$ $\left(\delta_{g}\mathcal{W}_{\mathscr{R}_{g}}+\delta_{l}\mathcal{W}_{\mathscr{R}_{l}}+\delta_{nl}\mathcal{W}_{\mathscr{R}_{nl}}+\mathcal{I}\right)^{-1}$
$\odot\Big{(}\delta_{g}\mathscr{R}_{g}^{\top}([\mathcal{Y}^{k+1}]_{g})+\delta_{l}\mathscr{R}_{l}^{\top}([\mathcal{Y}^{k+1}]_{l})$
$+\delta_{nl}\mathscr{R}_{nl}^{\top}([\mathcal{Y}^{k+1}]_{nl})+(\mathcal{D}-\mathcal{S}^{k+1})\Big{)};$
@@
11:  $(\mathcal{S}^{k},[\boldsymbol{X}_{i}^{k}]_{g},[\boldsymbol{X}_{i}^{l}]_{l},[\boldsymbol{X}_{i}^{k}]_{nl},[\mathcal{G}^{k}]_{g},[\mathcal{G}^{k}]_{l},[\mathcal{G}^{k}]_{nl},\mathcal{L}^{k})$
To enhance the efficiency and effectiveness of the MLTL2p method, we implement a two-phase scheme on the multi-scale low-rank model and the P-BCD method. In the first phase, we employ only the global and local low-rank regularization, which is sufficient for identifying stripes and dead lines. In particular, we use Algorithm 2 to solve model (32) with the parameter $\delta_{nl}$ set to zero. In the second phase, we utilize the nonlocal low-rank regularization along with the global and local low-rank regularization. As the estimated HSI from the first phase has already removed most of the noise, we can perform only one block matching to obtain a fixed nonlocal block extraction $\mathscr{R}_{nl}$, which can significantly reduce the computational cost. By applying Algorithm 2 to solve model (32) in this phase, the algorithm can refine additional details in the HSI, thereby improving the quality of the restored HSI.
Lastly, we comment on some situations in HSI denoising and destriping. If Poisson noise exists in HSIs, one may conduct the Anscombe transformation [30] and then apply the MLTL2p method. Also, if the HSIs are severely contaminated by the dead lines or stripes, one may apply the median method or the midway equalization method [39], respectively, for preprocessing.
@@ 6 Experimental Results
In this section, we conduct numerical experiments for removing mixed noise in HSIs. We summarize the comparison of methods for HSI denoising and destriping in Table 1. Among them, we select the methods that utilize tensor representation for competing. These include BM4D [29] with FGLR [37] as a preprocessing step to remove stripes, as well as LRTD [12], SNLRSF [6], LRTFL0 [45], and the data-driven method QRNN3D [43]. All numerical experiments are implemented in Matlab R2018a and executed on a personal desktop (Intel Core i7 9750H at 2.60 GHz with 16 GB RAM).
Table 1: Comparison summary of the methods for HSI denoising and destriping.
Methods Form Address Correlations? Stripe Low-rank
Global Local
Nonlocal
BM4D [29] Tensor No No Yes No Hard-thresholding
FGLR [37] Matrix Yes No No $\|\cdot\|_{*}$ Graph Laplacian
LRMR [52] Matrix No Yes No $\|\cdot\|_{1}$ Go Decomposition
DLR [51] Matrix Yes No No $\|\cdot\|_{*}$$\cdot\|_{*}$
SSLR-SSTV
[47]
Matrix No Yes No $\|\cdot\|_{*}$$\cdot\|_{*}$
LRTD [12] Tensor Yes No No $\|\cdot\|_{2,1}$ 3-D HOSVD
SNLRSF [6] Tensor Yes No Yes $\|\cdot\|_{1}$ Successive SVD
LRTFL0 [45] Tensor Yes No No $\|\cdot\|_{1}$ Block Term Decomp.
MLTL2p Tensor Yes Yes Yes $\|\cdot\|_{2,p}$ Indep. 3-D HOSVD